Condensed financial data of Monopoly Corporation appear below A cash dividend was declared and paid in full to stockholders during the year Required: Solve for the missing numbers. (Enter any deductions and cash outflows as a negative value.) MONOPOLY CORPORATION Comparative Balance Sheet December 31 Current Year Prior Year Assets 40,600 32,900 70,900 2,900 209,000 011 (44,500) 380,200 S311,800 S 66,410 S Cash Accounts receivable Inventories Prepaid rent Property, plant, and equipment Accumulated depreciation Total assets 3,490 237,500 (60,400) Liabilities and Stockholders Equity Accounts payable Accrued Liabilities Notes payable (long-term) Contributed capital Retained earnings Total liabilities and stockholders equity $39,800 S 34,900 12,900 159,000 29,500 75,500 380,200311,800 10,000 134,500 59,000 136,900
media%2F586%2F5861c732-6ff3-4bf4-88fb-8f
media%2F51a%2F51a43508-700b-4cad-8dc0-fe
Answer #1

Missing information is as calculated below:

1. Account Receivable: Change in account receivable is -4,900 and opening account receivable is 32,900.

So closing account receivable will be $37,800

2. Inventory:

Change in inventory is -24,500 and opening inventory is 70,900.

So closing inventory will be $95,400

3. Depreciation expense is $15,900 as shown in the cash flow statement.

4. Income taxes is (441,200-15,900-9,900-98,500-299,000) = 17,900

5. Prepaid Rent is (2,900-3,490 ) = 590

6. Accrued liabilities is 10,000-12,900 = -2,900

7. Net cash provided by operating activities is total of 72,300+15,900-4,900-24,500+590-2,900+4,900 = 61,390

8. Additional capital contributed is $29,500

9. Net cash provided by financing activities is 25,810-61,390+28,500 = -7,080

10. Payment of cash dividend is -7,080 = 29,500-24,500- Cash dividend

Cash dividend = 29,500-24,500+7,080

= 12,080